A cell count was performed on a serous fluid specimen. As the fluid was cloudy, a 1:5 dilution was prepared. The counts were: 41 and 43; 5 squares on each side were counted. What is the average cell count per L?
Answer #1

Counts on two sides= 41, 43.

5 squares on both side were counted. So average count per square, combining both sides is: (41+43)/10 = 84/10 = 8.4.

Each square has an volume of 0.1 microlitre. = 10^-7 L.

So the average cell count = Dilution * Average count per square * 10^7.

= 5* 8.4 * 10^7.

= 4.2* 10^8 cells/litre